OpenStudy (anonymous):

Factor completely the polynomial or indicate that the polynomial is not factorable using integers. x^2 - (y - 6) ^2

hero (hero):

The polynomial you have posted is a difference of squares: a^2 - b^2 = (a + b)(a - b) In this case, a = x and b = y - 6 so x^2 - (y - 6)^2 = (x + (y - 6))(x - (y - 6)) = (x + y - 6)(x - y + 6)

OpenStudy (anonymous):

how do you determine the positive and negatives?

hero (hero):

It's just using the formula and applying algebra